OpenStudy (anonymous):

what is the simpliflied form of x^4/y^3/x^5/y^7

OpenStudy (anonymous):

is this a compound fraction? start with \[\frac{x^4}{y^3}\times\frac{ y^7}{x^5}\]

OpenStudy (anonymous):

then subtract the exponents to get \[\frac{y^4}{x}\]

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

yw